# Clock Auctions Augmented with Unreliable Advice

### v1

@@ 3.1 Lower Bound
We now show that apart from achieve the “best of both worlds” in terms of asymptotic robustness and consistency guarantees, our learning-augmented clock auction also achieves a tight asymptotic dependence on the $\epsilon$ parameter. Specifically, we show that there exists a family of instances for which any $(1+\epsilon)$-consistent auction has robustness at least $\Omega(\frac{1}{\epsilon}\log n)$ (the proof is deferred to Appendix B).
@@ 3.2 The Error-Tolerant Auction
Next, we introduce the ErrorTolerantauction, which is an extension of Mechanism 2. This auction takes as input an error tolerance parameter $\overline{\eta}\geq 1$, referred to as the error-tolerance parameter, which can be chosen by the mechanism designer. The only modification from Mechanism 2 to ErrorTolerantis that in Line 2, the targeted revenue is adjusted from $R_{t}$ to $R_{t}/\overline{\eta}$. The mechanism ErrorTolerant, when the prediction error $\eta$ is at most $\overline{\eta}$, achieves an approximation of $(1+\epsilon)\eta$ to the optimal welfare. Additionally, it maintains an approximation of $\left(\frac{2000\overline{\eta}(1+\epsilon)}{9\epsilon}+1\right)H_{n}$ to the optimal welfare at all times. Note that if $\overline{\eta}=1$, the auction reduces to Mechanism 2, preserving the corresponding consistency and robustness guarantees. The formal description of ErrorTolerantand the proof of the following theorem are deferred to Appendix C.
@@ 4 Clock Auctions Guaranteeing a Stronger Notion of Consistency
In this section, we turn to the more demanding notion of consistency, which requires a good approximation of the predicted solution irrespective of the prediction quality. We show that for the problem on hand, achieving good consistency is much harder than achieving good consistency. In particular, there is a separation in terms of the trade-off one can achieve for these two different notions of consistency and robustness. Specifically, we show that if one wants $\alpha$-consistency for some constant $\alpha$, not only can one not achieve $O(\log n)$ robustness, but one must also suffer $\Omega\left(n^{\frac{1}{\alpha-1}}\log n\right)$ robustness. Our bound demonstrates that if one wants to achieve $\Theta(\log n)$ robustness, $\alpha$ also needs to be $\Theta(\log n)$. We also propose an algorithm that asymptotically matches the lower bound. Formally, we show that:
@@
is $\alpha$-consistent and $\beta$-robust.
Note that for $\alpha=2$, we have $\beta=\Theta(n\log n)$. As the consistency guarantee $\alpha$ approaches 1, the robustness grows exponentially (for $\alpha=4/3$, $\beta=\Theta(n^{3}\log n)$, and so on). This is in sharp contrast with the standard consistency notion, as we can achieve $1+\epsilon$ consistency with $O(\log n)$ robustness for any arbitrarily small $\epsilon$. See Figure 1 for a demonstration of the asymptotic trade-off the mechanism obtains for different sizes of instances.
Figure 1: The trade-off between robustness and consistency when $\beta=n^{1/(\alpha-a)}H_{n}$ for different values of $n$
The clock auction, named FollowTheBindingBenchmark, also alternates between uniformly raising the bids of the unpredicted and predicted bidders while maintaining a careful balance between the revenue from the predicted set and the welfare of the unpredicted set. However, to achieve consistency, it requires that in every round, the revenue collected from the predicted set be at least $\frac{1}{\alpha}$ of the rejected welfare of the unpredicted set plus the current revenue, i.e., the lower bound of the welfare of the predicted set. Unless this constraint is satisfied, the clock auction will continue raising the bids for the predicted set. Please see Mechanism 4 for a formal description.
@@
7      Run $\textsc{UniformPrice}(N\setminus\hat{\texttt{OPT}},\mathbf{p})$ until:(i) $\max_{F\in\mathcal{F}}\text{rev}(F\cap N,\mathbf{p})\geq\frac{\beta}{4H_{n}}\cdot R^{P}_{t-1}$; OR (ii) $\max_{F\in\mathcal{F}}v(F\setminus N)\geq\frac{\beta}{4}\cdot R^{P}_{t-1}$ and $\max_{F\in\mathcal{F}}\text{rev}(F\cap N,\mathbf{p})\geq\frac{\beta}{4H_{n}}\cdot R^{P}_{t-1}$
@@ 13     $t\leftarrow t+1$
As before, we assume that $\mathcal{F}$ is a set system where the predicted set is completely disjoint from all maximal unpredicted sets. This is without loss by Observation 1. We first prove the robustness guarantee. All the missing proofs can be found in Appendix D.
Robustness Argument
@@
We are now ready to prove the robustness guarantees of the mechanism 4. Similarly as in Section 3 we will analyze cases whether or not the subset of $\hat{\texttt{OPT}}$ is outputted differently. We first consider the case where a subset of $\hat{\texttt{OPT}}$ is outputted. The intuition is that we only serve a subset of predicted bidders if we do not satisfy either condition in Line 4 but we secured good revenue from the predicted set in the previous iteration of the while loop. We formalize this intuition in the following lemma.
@@
Consistency Argument
We now turn toward the more technically challenging consistency guarantee. Recall again that a mechanism achieves $\alpha$-consistency if the it achieves an $\alpha$-approximation to the value of the predicted set, irrespective of the prediction quality. To this end, we first resolve the simpler case where we serve some subset of predicted set. In this case the guarantee is achieved by design, we include a formal proof of this for completeness in Appendix D.
@@
The more difficult case involves ensuring $\alpha$-consistencygiven that we output a subset of some unpredicted set. Intuitively we want to argue that by defining a large enough $\beta$, the revenue we obtained from the best unpredicted set is enough to “cover” the maximum potential welfare the predicted set can contain. To this end, we begin with some individual bounds of agent’s value in the predicted set based on when they rejected the price offer.
@@
Fix a set $S$ and index bidders in $S$ in non-increasing value order. Two conditions hold:
@@
$v_{k-1}\geq\frac{(\alpha-1)k+1}{(\alpha-1)(k-1)}v_{k}.\qed$
@@
$p_{k-1}<\frac{(\alpha-1)k+1}{(\alpha-1)(k-1)}\cdot p_{k}.\qed$
@@ 4.1 Lower Bound
We now show that the guarantee that Mechanism 4 obtained is asymptotically tight.
@@
The robustness of any deterministic clock auction that is augmented with a prediction $\hat{\texttt{OPT}}$ and satisfies $\alpha$-consistencyfor $\alpha\in[1+\epsilon,H_{n}]$ is $\Omega(n^{1/(\alpha-1)}\log{n})$
@@
Consider a feasibility constraint $\mathcal{F}$ defined by two disjoint maximal feasible sets $F_{1}$ and $F_{2}$, i.e., $N=F_{1}\cup F_{2}$, and a set $F\subseteq N$ is feasible if and only if $F\subseteq F_{1}$ or $F\subseteq F_{2}$. Let the sizes of these sets be $|F_{1}|=k_{1}$ and $|F_{2}|=k_{2}$, and let $F_{2}$ be the set predicted to be optimal, i.e., $\hat{\texttt{OPT}}=F_{2}$.
@@
be a set of $k_{2}$ values that can be matched to the $k_{2}$ bidders in $F_{2}$, such that the highest value among them is $v_{1}=1$ and the rest of the sequence is decreasing such that for every $i\in\{2,3,\dots,k_{2}\}$:
$v_{i}=\frac{(\alpha-1)(i-1)+\delta}{(\alpha-1)i+1}\cdot v_{i-1},$
@@
where $\delta^{\prime}\to 0$ as $\delta\to 0$. One of the crucial properties of the values in $V_{2}$ is that based on the rate of growth and part 1 of Claim 1, for every $i\in\{1,2,\dots,k_{2}-1\}$ it satisfies:
$(\alpha-1)\cdot i\cdot v_{i}<\sum_{j=i+1}^{k_{2}}v_{j}.$ (6)
@@
Case 1 (the auction $\mathcal{M}$ accepts a subset of $F_{1}$): For this case, we show that there exists an instance (not in $\mathcal{I}$) for which the auction fails to achieve $\alpha$-consistency. To verify this, let $x$ be the number of bidders in $F_{1}$ that did not drop out and note that, given the harmonic structure of the values in $V_{1}$, the smallest value among them will be at most $1/x$. As we showed above, this means that the maximum price offered to any one of the winning bidders in $F_{1}$ was $p_{1}\leq 1/x$. Now, consider an alternative instance that is identical in terms of the values of all the bidders that dropped out, but has value $1/x$ for all of the bidders that won. Note that the outcome of the auction would be the same for this alternative instance, since the only difference is regarding values of bidders that are higher than the price offered to them, so the welfare achieved by $\mathcal{M}$ in this alternative instance would be $x\cdot\frac{1}{x}=1$. However, the social welfare of $\hat{\texttt{OPT}}=F_{2}$ is more than $\alpha$, which is implied by Inequality (6) for $i=1$, which violates the $\alpha$-consistency constraint.
@@
Case 2 (the auction $\mathcal{M}$ accepts a strict subset of $\hat{\texttt{OPT}}$): For this case, we once again show that there exists an instance (not in $\mathcal{I}$) for which the auction fails to achieve $\alpha$-consistency. To verify this, let $x\leq k_{2}-1$ be the number of bidder in $\hat{\texttt{OPT}}$ that did not drop out and note that, by the definition of the values in $V_{2}$, the smallest value among them is $v_{x}$. As we showed above, this means that the maximum price offered to any one of the winning bidders in $\hat{\texttt{OPT}}$ was $p_{2}\leq v_{x}$. We, once again, consider the alternative instance that is identical in terms of the values of all the bidders that dropped out, but has value $v_{x}$ for all the bidders in $\hat{\texttt{OPT}}$ that won. The outcome of the auction would not be affected by this change, so the resulting welfare would be $x\cdot v_{x}$ but, using Inequality (6) for $i=x$, this violates the $\alpha$-consistency constraint.
@@
Case 3 (the auction $\mathcal{M}$ accepts every bidder in $\hat{\texttt{OPT}}$): For this case, note that no price offered to a bidder in $\hat{\texttt{OPT}}$ exceeded $v_{k_{2}}$, otherwise the adversarial assignment of values would ensure that the corresponding bidder would have dropped out and $\mathcal{M}$ would not accept all of them. Now consider consider an alternative instance that is identical with respect to the values of the bidders in $F_{1}$, but the values of everyone in $F_{2}$ are equal to $v_{k_{2}}$. Note that the outcome of the auction would be the same in this alternative instance, since the only difference is regarding values of bidders that are higher than the price offered to them. Therefore, using Equation (5), we can conclude that the welfare achieved by the auction in this instance would be
@@
$k_{2}\cdot v_{k_{2}}=k_{2}\cdot\frac{\Gamma\left(1+\frac{\alpha}{\alpha-1}\right)(k_{2}-1)!}{\Gamma\left(k_{2}+\frac{\alpha}{\alpha-1}\right)}+\delta^{\prime}=\frac{\Gamma\left(1+\frac{\alpha}{\alpha-1}\right)k_{2}!}{\Gamma\left(k_{2}+\frac{\alpha}{\alpha-1}\right)}+\delta^{\prime},$
@@
$\beta=\Omega\left(\frac{H_{k_{1}}\Gamma\left(k_{2}+\frac{\alpha}{\alpha-1}\right)}{\Gamma\left(1+\frac{\alpha}{\alpha-1}\right)k_{2}!}\right)=\Omega\left(k_{2}^{1/(\alpha-1)}\log{k_{1}}\right),$
where the last equation is inferred by following the same steps as we did at the end of the proof of Lemma 14. If we let $k_{1}=n^{c}$ and $k_{2}=n-n^{c}$ for any constant $c>0$, this yields the claimed robustness lower bound. ∎

### v2

@@ 3.1 Lower Bound
We now show that apart from achieving the “best of both worlds” in terms of asymptotic robustness and consistency guarantees, our learning-augmented clock auction also achieves a tight asymptotic dependence on the $\epsilon$ parameter. Specifically, we show that there exists a family of instances for which any $(1+\epsilon)$-consistent auction has robustness at least $\Omega(\frac{1}{\epsilon}\log n)$ (the proof is deferred to Appendix B).
@@ 3.2 The Error-Tolerant Auction
This type of error is called the quality of recommendation by [CSV24]. Next, we introduce the ErrorTolerant auction, which is an extension of Mechanism 2. This auction takes as input a parameter $\overline{\eta}\geq 1$, referred to as the error-tolerance parameter, which is chosen by the mechanism designer. The only modification from Mechanism 2 to ErrorTolerant is that in Line 2, the targeted revenue is adjusted from $R_{t}$ to $R_{t}/\overline{\eta}$. When the prediction error $\eta$ is at most $\overline{\eta}$, ErrorTolerant achieves an approximation of $(1+\epsilon)\eta$ to the optimal welfare. Additionally, it maintains an approximation of $\left(\frac{2000\overline{\eta}(1+\epsilon)}{9\epsilon}+1\right)H_{n}$ to the optimal welfare at all times. Note that if $\overline{\eta}=1$, the auction reduces to Mechanism 2, preserving the corresponding consistency and robustness guarantees. The formal description of ErrorTolerant and the proof of the following theorem are deferred to Appendix C.
@@ 4 Clock Auctions Guaranteeing a Stronger Notion of Consistency
In this section, we turn to the more demanding notion of consistency, which requires a good approximation of the predicted solution irrespective of the prediction quality. We show that achieving good consistency is much harder than achieving good consistency. In particular, there is a separation in terms of the robustness that one can achieve for each of these two notions of consistency. Specifically, we show that if one wants $\alpha$-consistency for some constant $\alpha$, not only is it impossible to achieve $O(\log n)$ robustness, as we achieved for the standard notion of consistency in the previous section, but one must actually suffer $\Omega\left(n^{\frac{1}{\alpha-1}}\log n\right)$ robustness. Before proving this separation, we provide a learning-augmented clock auction that matches this bound for any constant $\alpha$.
@@
$\beta\geq\frac{8H_{n}\Gamma\left(n+\frac{\alpha}{\alpha-1}\right)}{\Gamma\left(1+\frac{\alpha}{\alpha-1}\right)n!}-\frac{4H_{n}(\alpha-1)}{\alpha}=\Theta\left(n^{\frac{1}{\alpha-1}}\log n\right),$
[…]
The clock auction defined in Mechanism 4, named FollowTheBindingBenchmark, also alternates between uniformly raising the bids of the unpredicted and predicted bidders while maintaining a careful balance between the revenue from the predicted set and the welfare of the unpredicted set. However, to achieve consistency, it requires that in every round the revenue collected from the predicted set be at least $\frac{1}{\alpha}$ of the rejected welfare of the unpredicted set plus the current revenue, i.e., the lower bound of the welfare of the predicted set. Unless this constraint is satisfied, the clock auction will continue raising the bids for the predicted set. See Figure 1 for a demonstration of the asymptotic trade-off the mechanism obtains for different sizes of instances.
Figure 1: The trade-off between robustness and consistency when $\beta=n^{1/(\alpha-1)}H_{n}$ for different values of $n$
@@
7      Run $\textsc{UniformPrice}(N\setminus\hat{\texttt{OPT}},\mathbf{p})$ until:(i) $\max_{F\in\mathcal{F}}\text{rev}(F\cap N,\mathbf{p})\geq\frac{\beta}{4}\cdot R^{P}_{t-1}$; OR (ii) $\max_{F\in\mathcal{F}}v(F\setminus N)\geq\frac{\beta}{4}\cdot R^{P}_{t-1}$ and $\max_{F\in\mathcal{F}}\text{rev}(F\cap N,\mathbf{p})\geq\frac{\beta}{4H_{n}}\cdot R^{P}_{t-1}$
@@ 13     $t\leftarrow t+1$
As before, we assume that $\mathcal{F}$ is a set system where the predicted set is completely disjoint from all maximal unpredicted sets; by Observation 1, this is without loss of generality. We first prove the robustness guarantee. All the missing proofs can be found in Appendix D.
Robustness Argument.
@@
We are now ready to prove the robustness guarantees of Mechanism 4. Similarly as in Section 3 we will analyze the two cases which depend on whether or not a subset of $\hat{\texttt{OPT}}$ is output separately. We first consider the case where a subset of $\hat{\texttt{OPT}}$ is output. The intuition is that we only serve a subset of predicted bidders if we do not satisfy either condition in Line 4 but we secured good revenue from the predicted set in the previous iteration of the while loop. We formalize this intuition in the following lemma.
@@
Consistency Argument.
We now turn toward the more technically challenging consistency guarantee. Recall again that a mechanism achieves $\alpha$-consistency if it achieves an $\alpha$-approximation to the value of the predicted set, irrespective of the prediction quality. To this end, we first resolve the simpler case where we serve some subset of predicted set. In this case the guarantee is achieved by design, but we include a formal proof of this for completeness in Appendix D.
@@
The more difficult case involves ensuring $\alpha$-consistency given that we output a subset of some unpredicted set. Intuitively we want to argue that by defining a large enough $\beta$, the revenue we obtained from the best unpredicted set is enough to “cover” the maximum potential welfare the predicted set can contain. To this end, we begin with some individual bounds of bidders’ values in the predicted set based on when they rejected the price offer.
@@
Fix a set of bidders $S$ and index bidders in $S$ in non-increasing value order. For all integers $1\leq k<|S|$, the following two conditions hold:
@@
$v_{k-1}\geq\frac{(\alpha-1)k+1}{(\alpha-1)(k-1)}v_{k},$
as desired. ∎
@@
$p_{k-1}<\frac{(\alpha-1)k+1}{(\alpha-1)(k-1)}\cdot p_{k},$
as desired. ∎
@@ 4.1 Lower Bound
We now provide a lower bound for all deterministic clock auctions in terms of the trade-off between consistency and robustness. This lower bound shows that the trade-off achieved by Mechanism 4 is near-optimal; in particular for any constant $\alpha$, the trade-off achieved by the mechanism is asymptotically optimal. We note that the lower bound below given in Theorem 5 is interesting only for $\alpha$ with $\alpha^{\alpha}=o(n)$.
@@
The robustness of any deterministic clock auction that is augmented with a prediction $\hat{\texttt{OPT}}$ and satisfies $\alpha$-consistencyfor $\alpha\in[1+\epsilon,H_{n}]$ is $\Omega(\frac{1}{\alpha}n^{1/(\alpha-1)}\log{n})$.
Note that for $\alpha=2$, we have $\beta=\Theta(n\log n)$. As the consistency guarantee $\alpha$ approaches 1, the robustness grows exponentially (e.g., for $\alpha=4/3$, $\beta=\Theta(n^{3}\log n)$). This is in sharp contrast with the standard consistency notion, for which we achieve $(1+\epsilon)$-consistency with $O(\log n)$-robustness for any arbitrarily small $\epsilon$.
Implications of the impossibility result on the feasible smoothness guarantees.
Before presenting the proof of this impossibility result, we would like to first point out its implications regarding the “price of smoothness,” i.e., the best achievable robustness if we require that the performance of the auction degrades smoothly as a function of the prediction error. Although a lot of prior work in the learning-augmented framework aims to design mechanisms with smoothness guarantees, to the best of our knowledge, (almost) none of this work provides explicit impossibility results regarding the price of smoothness. In contrast to this paucity of impossibility results, the robustness lower bounds that we prove for $\alpha$-$\text{consistent}^{\infty}$ auctions imply that any clock auction whose performance degrades linearly as a function of the error, $\eta$, needs to be $\Omega(\text{poly}(n))$-robust.
This bound on the price of smoothness is implied by the following general observation (which holds beyond our setting, as long as the error is measured as in (3), termed quality of recommendation by [CSV24]: any mechanism whose performance degrades linearly as a function of the error $\eta$, e.g., guaranteeing an approximation factor of $\alpha\eta$ if the prediction error is $\eta$, also satisfies $\alpha$-consistency. To verify this, note that the value of a predicted set $\hat{\texttt{OPT}}$ with error $\eta$ is $\text{OPT}/\eta$ (by definition). A mechanism with approximation factor $\alpha\eta$ guarantees welfare $\frac{\text{OPT}}{\alpha\eta}=\frac{\hat{\texttt{OPT}}}{\alpha}$, which implies a consistency of $\alpha$. This leads to the following Corollary regarding the price of smoothness.
Corollary 3.
If a deterministic clock auction augmented with a prediction $\hat{\texttt{OPT}}$ guarantees an approximation of $\alpha\eta$ for every $\alpha\in[1+\epsilon,H_{n}]$, where $\eta$ is the prediction error (as defined in (3)), then its robustness cannot be better than $\Omega(\frac{1}{\alpha}n^{1/(\alpha-1)}\log{n})$.
We now provide the proof of Theorem 5.
@@
(Proof of Theorem 5) Let $\alpha$ be the targeted consistency any mechanism is aiming for. Consider a feasibility constraint $\mathcal{F}$ defined by two disjoint maximal feasible sets $F_{1}$ and $F_{2}$, i.e., $N=F_{1}\cup F_{2}$, and a set $F\subseteq N$ is feasible if and only if $F\subseteq F_{1}$ or $F\subseteq F_{2}$. Let the sizes of these sets be $|F_{1}|=k_{1}$ and $|F_{2}|=\alpha k_{2}$, and let $F_{2}$ be the set predicted to be optimal, i.e., $\hat{\texttt{OPT}}=F_{2}$.
@@
be a set of $k_{2}$ values that can be assigned to the $\alpha k_{2}$ bidders in $F_{2}$. In particular, each value $v_{i}$ for $i\in[1,k_{2}-1]$ is assigned to a single bidder, while $v_{k_{2}}$ is assigned to the remaining $(\alpha-1)k_{2}+1$ bidders. The values in $V_{2}$ are in decreasing sequence such that the highest value among them is $v_{1}=1$ and, inductively, for every $i\in\{2,3,\dots,k_{2}\}$:
$v_{i}=\frac{(\alpha-1)(i-1)+\delta}{(\alpha-1)i+1}\cdot v_{i-1},$ (5)
@@
where $\delta^{\prime}\to 0$ as $\delta\to 0$.
By Equation (5), we have that $(\alpha-1)\cdot(k_{2}-1)\cdot v_{k_{2}-1}<((\alpha-1)k_{2}+1)v_{k_{2}}$. Note that the right-hand side is the sum of the values of the bidders in $F_{2}$ with value $v_{k_{2}}$. Again by Equation (5) and part 1 of Claim 1 we have, for every $i\in\{1,2,\dots,k_{2}-1\}$ it satisfies:
$(\alpha-1)\cdot i\cdot v_{i}<((\alpha-1)k_{2}+1)v_{k_{2}}+\sum_{j=i+1}^{k_{2}-1}v_{j}.$ (7)
@@
Case 1 (the auction $\mathcal{M}$ accepts a subset of $F_{1}$): For this case, we show that there exists an instance (not in $\mathcal{I}$) for which the auction fails to achieve $\alpha$-consistency. To verify this, let $x$ be the number of bidders in $F_{1}$ that did not drop out and note that, given the harmonic structure of the values in $V_{1}$, the smallest value among them will be at most $1/x$. As we showed above, this means that the maximum price offered to any one of the winning bidders in $F_{1}$ was $p_{1}\leq 1/x$. Now, consider an alternative instance that is identical in terms of the values of all the bidders that dropped out, but has value $1/x$ for all of the bidders that won. Note that the outcome of the auction would be the same for this alternative instance, since the only difference is regarding values of bidders that are higher than the price offered to them, so the welfare achieved by $\mathcal{M}$ in this alternative instance would be $x\cdot\frac{1}{x}=1$. However, the social welfare of $\hat{\texttt{OPT}}=F_{2}$ is more than $\alpha$, which is implied by Inequality (7) for $i=1$ since the left-hand side of $(\alpha-1)v_{1}=\alpha-1$ is less than the sum of all the remaining values in $F_{2}$, which violates the $\alpha$-consistency constraint.
@@
Case 2 (the auction $\mathcal{M}$ accepts a subset of $\hat{\texttt{OPT}}$ of fewer than $k_{2}$ bidders): By the fact that we may assign values adversarially to ensure that the first $(\alpha-1)k_{2}+1$ bidders whose clocks are raised above $v_{k_{2}}$ have value $v_{k_{2}}$, we may assume that all bidders with value $v_{k_{2}}$ exit the auction. For this case, we once again show that there exists an instance (not in $\mathcal{I}$) for which the auction fails to achieve $\alpha$-consistency. To verify this, let $x\leq k_{2}-1$ be the number of bidders in $\hat{\texttt{OPT}}$ that did not drop out and note that, by the definition of the values in $V_{2}$, the smallest value among them is $v_{x}$. As we showed above, this means that the maximum price offered to any one of the winning bidders in $\hat{\texttt{OPT}}$ was $p_{2}\leq v_{x}$. We, once again, consider the alternative instance that is identical in terms of the values of all the bidders that dropped out, but has value $v_{x}$ for all the bidders in $\hat{\texttt{OPT}}$ that won. The outcome of the auction would not be affected by this change, so the resulting welfare would be $x\cdot v_{x}$ but, using Inequality (7) for $i=x$ for any $x\leq k_{2}-1$, this violates the $\alpha$-consistency constraint since the left-hand side of Inequality (7) is $(\alpha-1)$ times the welfare the auction obtains and the right-hand side is the rejected welfare from the predicted set.
@@
Case 3 (the auction $\mathcal{M}$ accepts a subset of $\hat{\texttt{OPT}}$ of at least $k_{2}$ bidders): For this case, note that by the adversarial assignment of values, the auction maximizes its welfare if no price offered to a bidder in $\hat{\texttt{OPT}}$ exceeded $v_{k_{2}}$, otherwise the adversarial assignment would ensure that the corresponding bidder would have dropped out (and $\mathcal{M}$ then would not accept all bidders of value $v_{k_{2}}$). Now consider an alternative instance that is identical with respect to the values of the bidders in $F_{1}$, but the values of everyone in $F_{2}$ are equal to $v_{k_{2}}$. Note that the outcome of the auction would be the same in this alternative instance, since the only difference is regarding values of bidders that are higher than the price offered to them. Therefore, using Equation (6), we conclude that the welfare achieved by the auction in this instance would be
@@
$\alpha\cdot k_{2}\cdot v_{k_{2}}=\alpha\cdot k_{2}\cdot\frac{\Gamma\left(1+\frac{\alpha}{\alpha-1}\right)(k_{2}-1)!}{\Gamma\left(k_{2}+\frac{\alpha}{\alpha-1}\right)}+\delta^{\prime}=\frac{\alpha\Gamma\left(1+\frac{\alpha}{\alpha-1}\right)k_{2}!}{\Gamma\left(k_{2}+\frac{\alpha}{\alpha-1}\right)}+\delta^{\prime},$
@@
$\beta=\Omega\left(\frac{H_{k_{1}}\Gamma\left(k_{2}+\frac{\alpha}{\alpha-1}\right)}{\alpha\Gamma\left(1+\frac{\alpha}{\alpha-1}\right)k_{2}!}\right)=\Omega\left(\frac{1}{\alpha}k_{2}^{1/(\alpha-1)}\log{k_{1}}\right),$
where the last equation is inferred by following the same steps as we did at the end of the proof of Lemma 14. If we let $k_{1}=n^{c}$ and $k_{2}=\frac{n-n^{c}}{\alpha}$, this yields the claimed robustness lower bound. ∎